You decided to have a morning jog to keep your body fit. You jog 100 m East in 150 s, make a left turn and jog 150 m in 180 s, and finally make another left turn for 100 m in 90 s. 
1.    What is your average speed for the entire jog?
2.    What is your average velocity for the entire jog?
3.    How do you compare the magnitude of the average speed and velocity?

 
1
Expert's answer
2021-03-31T09:32:32-0400

1. Average speed (total distance over total time):


"s=\frac{100+150+100}{150+180+90}=0.833\text{ m/s}."

2. Average velocity is total displacement over total time:

"v=\frac{150}{150+180+90}=0.357\text{ m/s}."

3. Average speed is greater than average velocity.